[Hallc_running] [New Logentry] RC Report for Tuesday, March 3

brads at jlab.org brads at jlab.org
Tue Mar 3 20:45:02 EST 2020

Logentry Text:
<p><strong><a class="new" title="RC Meeting and Run Plan, Wednesday, March 4 2020" href="https://hallcweb.jlab.org/wiki/index.php?title=RC_Meeting_and_Run_Plan,_Wednesday,_March_4_2020">RC Meeting and Run Plan, Wednesday, March 4 2020</a></strong></p><p><strong>Tuesday 3 Mar 2020. </strong></p><h1><span id="General_updates" class="mw-headline">General updates</span></h1><ul><li>Target has been running smoothly at around 53%+, but Accelerator has been having problems<br><ul><li>Took ~3 hours of transverse data during Swing before beam was lost, then switched back to longitudinal (only managed ~15min before beam studies started)</li></ul></li></ul><ul><li>Injector did their spot move yesterday at 11am, beam back around 3pm</li><li>Rough beam through Swing and Owl<ul><li>Ran OK from 3pm--6pm, then Accel down until end of Swing while experts tried to solve interception problems in the Injector <a class="external autonumber" href="https://logbooks.jlab.org/entry/3796203" rel="nof!
 ollow">[1]</a><ul><li>Problems were not driven by the spot move, these are apparently the same interception issues that hit us a few days back.</li></ul></li><li>Had a few more hours of trippy beam in OWL until beam off again at 5am</li><li>More work on Injector and beam back around 8am for a little production before Beam Studies program that will run through Day shift</li></ul></li></ul><ul><li>Beam studies will attempt understand what is going on with the machine (beam loss, activation, etc) through Tuesday DAY shift <a class="external autonumber" href="https://opsweb.acc.jlab.org/CSUEApps/atlis/task/20368" rel="nofollow">[2]</a><ul><li>Test program starts around 9am with some beam into Hall C to study Hall C beam-correlated loss in the NL and east arc (in particular)</li><li>Target group entered around 10am and was out at 6pm, then MCC took another couple hours to restore beam and complete their test program</li><li>It sounds like the Beam Studies was productive and they!
  better understand some issues that have been plaguing the mac!
 hine. They wanted to think about the data a bit and will present some options Wednesday<ul><li>Many thanks to RadCon and OPs for sticking around a little longer while the Target folks worked to wrap up in the Hall</li></ul></li></ul></li></ul><ul><li>Target group + RadCon was in the Hall from around 10am -- 6pm with work focused on EPR issues<ul><li>They had a few issues, but managed to get to a working state by the time they had to get out</li><li>See <a class="external text" href="https://logbooks.jlab.org/entry/3796852" rel="nofollow">Hall Access for EPR Debugging and EPR NMR Calibration</a> for more details</li></ul></li></ul><ul><li>Mike Fowler rebooted the Magnet controls IOC in mid-Day shift to correct some issues<ul><li>GUI wasn't working quite right, not sure what else may have been affected</li></ul></li></ul><ul><li>Joe Did a quick walkthrough training for one person</li></ul><h2><span id="Notes_for_the_Future" class="mw-headline">Notes for the Future</span></h2>!
 <ul><li>'Macro Pulse Generator' replacement in injector will halt beam operations on Wed (8:30--1:30pm)<ul><li>Hall C may take another opportunistic access in that period (TBD)</li></ul></li></ul><ul><li>Joe has two other folks that need a Hall walk through during an opportunistic down<ul><li>Could happen Wed between 10--noon, if they respond</li></ul></li></ul><ul><li>Schedule <strong>Moller</strong> measurement for <strong>Thursday DAY</strong> shift</li></ul><ul><li>'Filler' runs for next spin up / target-down period:<ul><li>1 hour run with target out for measure of the background just from the beryllium windows</li></ul></li></ul><h2><span id="Pending_Issues" class="mw-headline">Pending Issues</span></h2><ul><li>pNMR issues (not sure this was looked at yet?)</li><li>EPR problems <a class="external autonumber" href="https://logbooks.jlab.org/entry/3793821" rel="nofollow">[3]</a> saw some progress on Tuesday <a class="external autonumber" href="https://logbooks.jlab.org/e!
 ntry/3796852" rel="nofollow">[4]</a>, but issues remain<ul><li>Need a n!
 ew fiber bundle ASAP at minimum</li></ul></li><li>Kepco power supply commissioning/installation (Todd, Bill)<ul><li>Bill will update status on Wed.</li></ul></li><li>Verify that calibration constants and input values associated with NMR and other polarimetry are consistent and will be correct when the software is restarted <a class="external autonumber" href="https://logbooks.jlab.org/entry/3796178" rel="nofollow">[5]</a> <a class="external autonumber" href="https://logbooks.jlab.org/entry/3795421" rel="nofollow">[6]</a> <a class="external autonumber" href="https://logbooks.jlab.org/entry/3795982" rel="nofollow">[7]</a> <a class="external autonumber" href="https://logbooks.jlab.org/entry/3796117" rel="nofollow">[8]</a> <a class="external autonumber" href="https://logbooks.jlab.org/entry/3796111" rel="nofollow">[9]</a> (Bill, Junhao)</li><li>Persistent BLM & ion chamber trips<ul><li>General activation, weird BLM issues (<a class="external free" href="https://logbooks.jla!
 b.org/entry/3792326" rel="nofollow">https://logbooks.jlab.org/entry/3792326</a>), Ad-hoc ion chamber threshold adjustments (<a class="external free" href="https://logbooks.jlab.org/entry/3792095" rel="nofollow">https://logbooks.jlab.org/entry/3792095</a>)</li><li><strong>NOTE</strong>: Shift crew should <em>NOT</em> give authorization to MCC change ion chamber thresholds. If they ask, tell Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)</li><li>Beam Studies program will investigate some of these issues (most pressing is significant activation + beam loss in the arcs)</li></ul></li><li>Contact with Target readbacks through AnywhereUSB USB bridge still drop out on occasion<ul><li>Initial replacement of the USB bridge failed (did not seem to help underlying issue, then bridge stopped working), replaced with original device</li></ul></li><li>OPs wished to re!
 boot hung iochc10 <a class="external autonumber" href="https://logbooks!
 .jlab.org/entry/3795636" rel="nofollow">[10]</a>, noted in December as well <a class="external autonumber" href="https://logbooks.jlab.org/entry/3745818" rel="nofollow">[11]</a><ul><li>This runs the Moller cryo controls and should not be rebooted without expert oversight (does not impact operations). D. Gaskell is looking into getting this addressed properly with OPs.</li></ul></li></ul><h2><span id="Opportunistic_Access_Jobs" class="mw-headline">Opportunistic Access Jobs</span></h2><h2><span id="Target_coil_rotation_planned_for_March_13--20" class="mw-headline"><strong>Target coil rotation planned for March 13--20</strong></span></h2><ul><li>Nominal 1-week downtime in Hall C to rotate target coils, install new cell, and transition to d2n program</li></ul><ul><li>More details are here: <a title="PolHe3 Target Coil Rotation Plan, Mar 2020" href="https://hallcweb.jlab.org/wiki/index.php/PolHe3_Target_Coil_Rotation_Plan,_Mar_2020">PolHe3_Target_Coil_Rotation_Plan,_Mar_2020</a>!


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3796870
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200303/4324826a/attachment.html>

More information about the Hallc_running mailing list